OpenStudy (anonymous):

A single electron orbits a lithium nucleus that contains three protons (+3e). The radius of the orbit is 1.76E-11m. Determine the kinetic energy of the electron.

OpenStudy (anonymous):

Is this Intro Physics E&M or Modern Physics+ ?

OpenStudy (anonymous):

I'm not sure what either of those mean. This is college physics 2

OpenStudy (anonymous):

Then it is intro.. I ask that so that if I need to look up or do any other things.. But since this is intro, we only have to worry about Orbit Radius, Coulomb Force and momentum to come up with a number in Joules. Which is where we start..

OpenStudy (anonymous):

Well coulumb's law is what most of this chapter is about. i looked up physics e&m and i suppose this would fall under that category

OpenStudy (anonymous):

I'm just not sure how to apply it to KE

OpenStudy (anonymous):

Ok.. I will work you though it.. I was curious what Carlos was writing.. but I might as well start.. First get is the mass of the electron.. You can treat it as if it was a planet moving around the sun.. In that case, your Coulomb force is your "gravity" with sign convention that can basicly be ignored.

OpenStudy (anonymous):

Ok, 9.11E-31

OpenStudy (anonymous):

should i figure out coulomb force?

OpenStudy (anonymous):

Yeah. Go a head and calculate that up..

OpenStudy (anonymous):

Centrifugal force must equal electrostatic attraction:\[F_c=m_e·\frac{ v^2 }{ r }\]\[F_E=\frac{ 1 }{ 4 \pi \epsilon_0 }\frac{ 3e^2 }{ r^2 }\] \[F_E=F_C \rightarrow m_e \frac{ v^2 }{ r }=\frac{ 1 }{ 4 \pi \epsilon_0 }\frac{ 3e^2 }{ r^2}\rightarrow E_K=\frac{ 1 }{ 2 } m_ev^2=\frac{ 3e^2 }{ 8 \pi \epsilon_0r}\]

OpenStudy (anonymous):

can you follow that?

OpenStudy (anonymous):

Not really

OpenStudy (anonymous):

thank you though carlos

OpenStudy (anonymous):

He basicly took 2 Equations and substituted them together.

OpenStudy (anonymous):

welcome. what is it that you dont understand?

OpenStudy (anonymous):

well is Me mass of the electron?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

why are you only counting the proton charges and not the electron charges?

OpenStudy (anonymous):

Well.. How much is the charge of an electron?

OpenStudy (anonymous):

1e=1.6E-19 correct?

OpenStudy (anonymous):

why is that negated?

OpenStudy (anonymous):

which is the same as a proton.. Just with a flipped charge sign.

OpenStudy (anonymous):

right, there are 4 charges in total. -e and +3e why is the -e dropped?

OpenStudy (anonymous):

I am counting both. Coulomb force between q1 and q2 is:\[F_E=\frac{ 1 }{ 4 \pi \epsilon_0 }\frac{ q_1q_2 }{ r^2 }\]Now see that q1=3e (three protons) and q2=e (one electron) whwre e=1.6·10^(-19) coulombs and you get my formula

OpenStudy (anonymous):

na.. his is right. (sorry thought you were saying something else, my fault)

OpenStudy (anonymous):

yeah but up there it goes from (3e)(e)/r^2. why is the q2 dropped?

OpenStudy (anonymous):

ohhhhh

OpenStudy (anonymous):

If Earth is orbiting around the sun.. the mass of the earth is multiplied into the mass of the sun... not added... That's where you are confused.

OpenStudy (anonymous):

nothing dropped, the formula says e^2

OpenStudy (anonymous):

yeah it just clicked

OpenStudy (anonymous):

got it now?

OpenStudy (anonymous):

yeah, i do actually just had to go step by step in that formula on paper

OpenStudy (anonymous):

thanks a lot for being so patient guys

OpenStudy (anonymous):

awesome.. np.. I used to be a tutor for my school. :)

OpenStudy (anonymous):

awesome, you guys are good. thanks again i really appreciate it

OpenStudy (anonymous):

anyway.. what is the next thing sticking you?

OpenStudy (anonymous):

By the way, this is the Rutherford model for the atom, which is no longer valid and was replaced by Bohr´s model

OpenStudy (anonymous):

well that I will have to see. I have been sitting on this problem for hours. I know ill run into more speed bumps. mind if i message you later if youre on?

OpenStudy (anonymous):

the bohr is the cloud model?

OpenStudy (anonymous):

yep, based on probability

OpenStudy (anonymous):

the Ruth model was basicly saying atoms travel in descrete orbits.. like planets.. which is all you end up doing in intro phy 2.

OpenStudy (anonymous):

ok, i feel this is merely a stepping stone. bohr would complicate things further

OpenStudy (anonymous):

atoms... pshh.,.. I mean electrons orbit

OpenStudy (anonymous):

i got what you were saying

OpenStudy (anonymous):

So, back to it.. Are you seeing the Kinetic Energy from Carlos' last line?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

KE=coulombs force

OpenStudy (anonymous):

sort of i mean

OpenStudy (anonymous):

:( Force is not energy...

OpenStudy (anonymous):

I see the math, the understanding of it all is lost. However I think if i stick with it knowing the solution i will figure it out

OpenStudy (anonymous):

"Work" is energy. You have to find out how much energy there is by Work done by the force on the object in question.

OpenStudy (anonymous):

I move from centrifugal forces to Kinetic Energy multiplying by 1/2r and do the same with Coulomb force

OpenStudy (anonymous):

I urge you not to use the solution as a means of understanding... (just saying)

OpenStudy (anonymous):

yeah, you put it better. Working backwards is my best ally when it comes to understanding concepts.

OpenStudy (anonymous):

@QuantumQQ I concur

OpenStudy (anonymous):

well guys, I have to go. @scottman you are in good hands, bye

OpenStudy (anonymous):

ok thanks guys i appreciate it.

OpenStudy (anonymous):

thanks @CarlosGP

OpenStudy (anonymous):

actually I found that really the directions to be the hardest part.. It is also easy to make up interpretations.. Up as Carlos's equations are put.. and with the question... You want to make the Fe = FE <------ important key (take care carlos)

OpenStudy (anonymous):

Fe meaning force of electron and FE meaning electric field correct?

OpenStudy (anonymous):

And I will say.. Scroll up and look at the equations.. tell me what the KE equation is...

OpenStudy (anonymous):

1/2MeV^2

OpenStudy (anonymous):

Electrostatic <-- static meaning not moving, and electro meaning.. :)

OpenStudy (anonymous):

yup.. now.. tell me.. what are your unknowns in the KE equation?

OpenStudy (anonymous):

velocity

OpenStudy (anonymous):

yup.. and from the equations given.. where can we dig up a velocity?

OpenStudy (anonymous):

coulombs law?

OpenStudy (anonymous):

wait so if we get Force out of coulombs law and we know F=ma we know coulombs law=Me(a)

OpenStudy (anonymous):

no.. there is a v in the Fe Since, the Fe needs to be equal to the FE.. this would imply that orbit (r) is fixed and not changing (orbit collapsing or going out.)

OpenStudy (anonymous):

yeah... there is a little identity that says a = v^2/r

OpenStudy (anonymous):

(for rotational motion only that is)

OpenStudy (anonymous):

yeah so since we now would know r and a we just need to square root to find v?

OpenStudy (anonymous):

yeah.. write the equality in terms of the velocity squared, since there is no reason to unsquare it.... we can just directly insert the rewritten equation directly into 1/2 m_e v^2

OpenStudy (anonymous):

so then we have our asnwer?

OpenStudy (anonymous):

Yup. 1/2 m_e v^2 gives you the kinetic energy of the system.. We don't have to worry about any potential energies at all.. yay!

OpenStudy (anonymous):

oh wow, you made me conceptually understand this problem. You truly are awesome.

OpenStudy (anonymous):

Just tell me if the energy is positive or neg..... and you should be home free.

OpenStudy (anonymous):

positive

OpenStudy (anonymous):

awesome... nailed it. :)

OpenStudy (anonymous):

thank you. you have a gift. are you a teacher or prof of some sort?

OpenStudy (anonymous):

i know you said tutor but.. wow

OpenStudy (anonymous):

remember that if you are having trouble with a problem.. Chances are, the problem is really in the conceptual understanding... No.. I'm a tutor.. a Physics major.. doing my thing :) Thank you for the awesome compliment though :D

OpenStudy (anonymous):

Yeah sometimes it's hard bridging the gap without some assistance. Again thank you, would you mind if i messaged you if I run into any future snags?

OpenStudy (anonymous):

sure... I infrequently come here but I will try to get back to you if I get the message.

OpenStudy (anonymous):

Cool thanks again.

OpenStudy (anonymous):

(that's why I only have 53 as a score) Good luck with your class and take care :)

OpenStudy (anonymous):

you too